• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name fansler

    Log entry time 14:34:07 on September16,2005

    Entry number 154876

    keyword=Right Q1 status

    Began troubleshooting Right Q1. Magnet would not ramp above 86 amps. Checked all control and interlock systems, and discovered a bad He level meter. Replaced He level meter. Meter failed appoximately 10pm the evening prior. The valves closed, causing the magnet to warm to 115k, and spoil the vacuum. Placed a blower on the RQ1 insulating vacuum. Began cooldown approximately 10 am. Estimate cooldown completion approximately 3pm. We will then check out the system and resume troubleshooting if necessary.